OpenStudy (anonymous):

put y= -x/2 into y=mx+b form

jimthompson5910 (jim_thompson5910):

y= -x/2 y= -(1/2)x+0 The equation is now in y = mx+b form where m = -1/2 and b = 0

OpenStudy (anonymous):

how di du get 2 that step?

jimthompson5910 (jim_thompson5910):

-x/2 = -1/2x which is the same as -x/2 = -(1/2)x Now just add 0 to both sides -x/2+0 = -(1/2)x+0

OpenStudy (anonymous):

thnx sooo much u have been so helpful

jimthompson5910 (jim_thompson5910):

np